OpenStudy (anonymous):

A room contains 96 people. A man leaves and is replaced by a woman, leaving three times as many women as men in the room. How many men and how many women were there originally?

OpenStudy (anonymous):

Let's use 'M' and 'W' to represent the number of men and women in the room originally. Can you write an equation that states the total number of people was 96?

OpenStudy (anonymous):

i guess some thing like this, 96-M=96+W. but im not sure

OpenStudy (anonymous):

You're over-thinking it :-) ... it's just M + W = 96.

OpenStudy (anonymous):

So whats going to be the other equation?

OpenStudy (anonymous):

OK, let's call that equation 1 and leave it for a second. Now we need equation 2. First, how many men are there in the room (in terms of M) after one man leaves? And, how many women after a woman arrives?

OpenStudy (anonymous):

After one man leaves and women arrives

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

so that brings to this 96-M=96+W right?

OpenStudy (anonymous):

no. I'm thinking of something simpler. If there are M men in the room, and one leaves, now how many are there?

OpenStudy (anonymous):

M-1?

OpenStudy (anonymous):

Right. And after one woman arrives how many are there?

OpenStudy (anonymous):

W+1?

OpenStudy (anonymous):

Right. You're on your way. Now, let's say hypothetically, the question said that that after a man leaves, and a woman arrives, there are the same number of men and women in the room. How could you write that as an equation?

OpenStudy (anonymous):

M-1=W+1 ?

OpenStudy (anonymous):

yes, exactly! One more step. What the problem actually says is that, after the replacement, there are 3 times as many women in the room. Another way of saying this is that the number of women is equal to 3 times the number of men. Can you now translate that into an equation?

OpenStudy (anonymous):

\( 3\times (M-1) =( W+1) \)

OpenStudy (anonymous):

So is that the second equation?

OpenStudy (anonymous):

I think so.

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

Now solve it tot get M, and W ...

OpenStudy (anonymous):

Now you have 2 equations with 2 unknowns, so solve simultaneously. The substitution method is probably easiest here.

OpenStudy (anonymous):

Ok